# Causal Concept Embedding Models: Beyond Causal Opacity in Deep Learning

### v1

@@ Appendix A Architecture
Causal CEM provides versatile initialisation options for the adjacency matrix $A$, tailored to specific scenarios. In certain instances, weights can be derived from domain expertise or provided along with training data and labels. Without such information, weights can be directly inferred from training labels through causal structural learning algorithms [7] as a preliminary step. These approaches guide the model towards a predetermined decision-making pathway. Alternatively, weights can be learnt concurrently during the Causal CEM training phase, as outlined in Section 3.2. These weights may be initialised either randomly or based on the conditional entropy between labels, providing a better starting point. Additionally, a hybrid approach is feasible, where certain elements in $A$ are fixed while others remain trainable. For instance, a causal structural learning algorithm might yield a Partial Ancestral Graph (PAG) with undirected edges, allowing for the definition of directed edges and learning the direction for others to avoid cycle formation.
@@
In Causal CEM, the function $f_{i}$ corresponds to a causal mechanism in a SCM. Such mechanisms are typically formalized via structural equations. For instance, linear models are a common choice for label predictors in Concept Bottleneck Models [23] where endogenous embeddings are aggregated using a permutation invariant aggregator function $\oplus$ (such as the element-wise maximum, or sum):
$\hat{v}_{i}=\sigma\left(W_{i}\bigoplus_{j\in\{1,\dots,k\}}m_{ij}\hat{\mathbf{v}}_{j}+\mathbf{b}_{i}\right)$ (15)
However, other options are also available to increase the expressiveness and interpretability of the decision-making process, such as Deep Concept Reasoning [33] class predictors, which build logic-based formulae to obtain class label predictions using endogenous embeddings:
$\hat{v}_{i}\leftarrow\bigvee_{\mathbf{x}\in X_{\text{train}}}\bigwedge_{j\in\{1,\dots,k\}}l_{j}(\mathbf{x})=\bigvee_{\mathbf{x}\in X_{\text{train}}}\bigwedge_{j\in\{1,\dots,k\}}(\rho_{ij}(m_{ij}\hat{\mathbf{v}}_{j})\iff\hat{v}_{j}^{\prime})$ (16)
@@
Compositional generalization
The training procedure of Causal CEMs is highly parallelizable and modular as only direct connections need to be trained together (e.g., $a\rightarrow b$ and $b\rightarrow c$), while the model takes care of distant connections in an indirect way. For instance, the connection $a\rightarrow b\rightarrow c$ can be obtained as a composition of two different independent training procedures for $a\rightarrow b$ and $b\rightarrow c$. As a result, it is trivial for a Causal CEM to make the causal graph grow even at test time (see Figure 6). This can be done by composing two different graphs obtained by independent training procedures, encoders, datasets, or data types. We note that this is not possible in standard CBMs, which need to re-train the task predictor from scratch whenever new concepts or tasks are added to the mix. Moreover, this modularity enables a form of out-of-distribution compositional generalization as it creates new distant connections between variables that were never part of the same training procedure (e.g., $a$ and $c$ in the previous example).
Figure 6: Compositional generalization in Causal CEMs: two different Causal CEMs architectures are trained independently and then composed only at test time, thus creating a larger graph and allowing out-of-distribution causal inference.
Appendix B Experimental setup
B.1 Datasets
@@
  • dSprites [46] — The dataset comprises images featuring one of three objects (square, heart) in various positions and sizes. The defined concepts include: (1) object shape (square or heart), (2) object size (small or large), (3) vertical position (top or bottom), (4) horizontal position (left or right), (5) object colour (red or blue). Based on these, causal relationships and a binary classification task are established: if the object is a heart on the right side, it is large; if a heart is at the top of the image, it is red; the label is positive if the object is both red and large.
  • CelebA [47] — The CelebA dataset features celebrity images annotated with various attributes, including lipstick presence, gender, facial shape, and hair type. Gender is used as the classification label. This dataset is chosen for the presence of correlations and biases, such as the association between wearing lipstick and being identified as female.
B.2 Baselines
@@
  • Concept Bottleneck Model (CBM) [23]: This model first uses a concept predictor to infer concepts from the initial input, followed by a task label prediction based on these concepts. It is designed to be interpretable and treats concepts as significant informational to predict the task label.
  • Concept Embedding Model (CEM) [36]: Comprising $n$ context encoders, one for each concept, this model predicts each concept based on its respective context before predicting the final task label. It treats concepts and task labels in the same way as CBM.
B.3 Experiments
In our experiments, we evaluate our approach by examining four key dimensions: (1) performance accuracy, (ii) influence of ground-truth interventions, (iii) identification of causal structures, and (iv) blocking for the influence of one variable on another.
@@ B.3 Experiments
In the second aspect, we evaluate our approach by comparing it against CBM and CEM in terms of response to ground-truth interventions. Enhancing the impact of interventions in the Concept-Based Model is crucial for improving the role of humans in the loop. In our experiments, we initially perturbed the inputs to reduce label prediction accuracy, following methodologies established in prior research [36]. Subsequently, we implemented interventions on the most inaccurately predicted concepts in CEM and CBM. This intervention strategy is considered highly effective, as noted in [48]. For the Causal CEM, interventions began with concepts that have a higher number of descendant nodes in the model’s graph, aiming to maximise the intervention’s effectiveness. To assess this dimension, we measured the change in accuracy for non-intervened labels before and after the interventions on $n$ concepts (Delta Label Accuracy).
In the third aspect, we visualise the DAG utilised during the inference stage by Casual CEM and derive the corresponding logic equations. We generate Sum of Product logic rules from a table that lists all possible combinations of input concept values alongside the most frequent prediction for each combination derived from the training set, similarly to what done by [49]. It is crucial to note that while these logic rules are general for the model decision-making process, exogenous information may alter predictions for particular instances.
In the final dimension of our analysis, we compare Causal CEM, which operates on a specified graph, against CBM and CEM in terms of their efficacy in mitigating a variable’s influence on the task. Specifically, we perturb the prediction of a concept and then, following the graph structure utilised by Causal CEM, we intervene with the ground truth label on all descendant nodes (blocking). Consequently, in Causal CEM, all links between the altered concepts and the labels are deleted, a feat unachievable in CBM and CEM. To assess this characteristic, we calculated the Residual Concept Causal Effect, ratio of the Concept Causal Effect [44] post- and pre-application of the blocking techniques. Ideally, this ratio should be zero, indicating that after blocking, the altered node’s value no longer influences the outcome of the task.
B.4 Implementation details
Additional details To maximise the efficacy of interventions in Causal CEM, the second term of the loss can be regularised to maximise the average Causal Concept Effect (CaCE) [44] as follows:
@@ B.4 Implementation details
Here, $r$ represents a randomly chosen index for each sample, which is used to select one of the concepts following Espinosa Zarlenga et al. [36]. This regularisation can be weighted using an hyperparameter, $\lambda_{3}$. Moreover, for all the experiments where the graph is learnt end-to-end, we initialise the learnable adjacency matrix with the conditional entropy between each pair of values, extracted from the training set.
@@
Hyperparameters All baseline and proposed models were trained for varying epochs across different datasets: 500 for Checkmark, 200 for dSprites, and 30 for CelebA. The optimal epoch for each was determined based on label accuracy on the validation set. A uniform learning rate of 0.01 was applied across all models and datasets. For the CBM and CEM models, both concept and task losses were equally weighted at 1. This weighting scheme was also applied to the loss terms for endogenous copies’ prediction, endogenous variables’ prediction ($\lambda_{1}$), and graph priors ($\lambda_{2}$). The weight assigned to the loss terms in our models to maximize CaCE is 0.05. Additionally, $\gamma$ was treated as a learnable parameter, initialized at 0.1, and $\beta$ was set to 1. All experiments were conducted using five different seeds (1, 2, 3, 4, 5).
@@
Code, licenses and hardware For our experiments, we implement all baselines and methods in Python 3.9 and relied upon open-source libraries such as PyTorch 2.0 [50] (BSD license), PytorchLightning v2.1.2 (Apache Licence 2.0), Sklearn 1.2 [51] (BSD license). In addition, we used Matplotlib [52] 3.7 (BSD license) to produce the plots shown in this paper. Two datasets we used are freely available on the web with licenses: dSprites (Apache 2.0) and CelebA, which is released for non-commercial use research purposes only. We also introduce the Checkmark dataset and we described it in this section. We will publicly release the code with all the details used to reproduce all the experiments under an MIT license. The experiments were performed on a device equipped with an M3 Max and 36GB of RAM, without the use of a GPU. Approximately 40 hours of computational time were utilized from the start of the project, whereas reproducing the experiments detailed here requires only 2 hours.
Appendix C Additional results
@@
In this section, we include all the experiments shown in Section 5 for the three datasets in more detail and an ablation study on the value of $\lambda_{3}$.
C.1 Ground-Truth interventions
@@
Figure 7 illustrates the performance comparison among Causal CEM (using both the provided and learnt graphs), CBM, and CEM regarding the effects of interventions. Delta Label Accuracy, which quantifies the change in label accuracy before and after interventions on a growing number of concepts, is calculated solely for the concepts not directly intervened upon. In particular, Causal CEM demonstrates superior performance when interventions involve fewer concepts. This superior performance is attributed to the propagation of intervention effects through all descendant nodes in Causal CEM, unlike CBM and CEM, where the impact is confined to the final task without adjustments to other concepts. The most significant performance gain, approximately 15 percentage points, is observed in CelebA after seven interventions. This effect is particularly pronounced in scenarios with multiple concepts, such as in dSprites and CelebA. However, in simpler tasks with fewer concepts, like Checkmark, the advantage offered by our method is lower. Furthermore, when focusing solely on the effects of interventions on the task label, the causal graph utilised by Causal CEM allows us to identify beforehand the specific subset of concepts influencing the task prediction. This pre-identification significantly decreases the required number of interventions to achieve the desired outcome. Figure 8 demonstrates that Causal CEM attains comparable improvements in task performance but after interventions on only three or four concepts, in contrast to the ten and eleven concepts required by CEM and CBM, respectively. The elevated standard error observed in Causal CEM with the learnt graph is attributed to the variability of the graph structure, which significantly influences the outcomes of interventions.
C.2 Causal structures
@@
Figure 9 illustrates the adjacency matrices corresponding to the DAGs used by Causal CEM for inference in three datasets. On the other hand, Tables 3, 4, and 5 present the logic rules derived from the adjacency matrices depicted in the aforementioned figure. Notably, in the Checkmark dataset, both configurations successfully identified the ground truth graph and the correct logic rules. In the case of dSprites, the DAG identified through causal structural learning (GRaSP [40]) accurately discovers the causal graph and associated logic rules. Although the end-to-end model accurately identifies the correct relationships between concepts and tasks, it proposes alternative methods for concept prediction. It is important to note that even though the model did not identify the correct causal graph, the model was still capable of performing causal inference with the existing graph. In the CelebA data set, where there is no ground truth for either the graph or logic rules, the findings by GRaSP and the end-to-end model appear plausible and reveal biases inherent in the dataset, such as the strong correlation between makeup use and gender or potential causal links like smiling and a slightly open mouth. This scenario underscores the benefits of employing Causal CEM, particularly in demonstrating how specific concepts are used to predict other concepts and tasks.
C.3 Ablation study
@@
In Tables 6, 7, and 8, we present the outcomes of varying the hyperparameter $\lambda_{3}$, which weights the loss term designed to enhance the CaCE effect. The results indicate that optimising this loss term contributes to improved CaCE scores, thereby augmenting the efficacy of the interventions. Nonetheless, excessively high values of $\lambda_{3}$ may lead to diminished model performance, as it tends to prioritise boosting the CaCE score at the expense of accurate predictions.
@@
Figure 7: Impact of ground-truth interventions on concepts across three datasets. This figure illustrates the variations in accuracy for non-intervened labels, comparing performance before and after interventions on specific nodes.
@@
Figure 8: Impact on the task accuracy of ground-truth interventions performed on CelebA concepts. Causal CEM+CD has received in input a causal graph, discovered with a causal structural learning algorithm (GRaSP [40]), while Causal CEM learns it end-to-end.
@@
Figure 9: Adjaciency matrices representing the DAG used by Causal CEM during inference on the three datasets. On the left side, the matrices represent the DAG learnt end-to-end by the model, while on the right the DAG discovered with GRaSP [40]. It provides a PAG starting from the training data.
@@
Table 3: Logic rules extracted for the Checkmark dataset from Causal CEM+CD with a given DAG and from Causal CEM with a learnt DAG. A term which refers to an exogenous variable is omitted for simplicity.
@@
Causal CEM $\text{a}&\leftarrow\epsilon_{0}\\ \text{b}&\leftarrow\epsilon_{1}\\ \text{c}&\leftarrow\sim\text{b}\\ \text{d}&\leftarrow\text{a}\wedge\text{c}\\ \end{aligned}$
Causal CEM+CD $\text{a}&\leftarrow\epsilon_{0}\\ \text{b}&\leftarrow\epsilon_{1}\\ \text{c}&\leftarrow\sim\text{b}\\ \text{d}&\leftarrow\text{a}\wedge\text{c}\\ \end{aligned}$
Table 4: Logic rules extracted for the Dsprites dataset from Causal CEM+CD with a given DAG and from Causal CEM with a learnt DAG. A term which refers to an exogenous variable is omitted for simplicity.
@@
Causal CEM $\text{Shape}&\leftarrow\text{Size}\\ \text{Size}&\leftarrow\epsilon_{1}\\ \text{PosY}&\leftarrow\epsilon_{2}\\ \text{PosX}&\leftarrow\epsilon_{3}\\ \text{Color}&\leftarrow\text{Shape}\\ \text{Label}&\leftarrow\text{Size}\wedge\text{Color}\\ \end{aligned}$
Causal CEM+CD $\text{Shape}&\leftarrow\epsilon_{0}\\ \text{Size}&\leftarrow\text{Shape}\wedge\text{PosX}\\ \text{PosY}&\leftarrow\epsilon_{2}\\ \text{PosX}&\leftarrow\epsilon_{3}\\ \text{Color}&\leftarrow\text{Shape}\wedge\text{PosY}\\ \text{Label}&\leftarrow\text{Size}\wedge\text{Color}\\ \end{aligned}$
Table 5: Logic rules extracted for the Celeba dataset from Causal CEM+CD with a given DAG and from Causal CEM with a learnt DAG. A term which refers to an exogenous variable is omitted for simplicity.

### v4

@@ Appendix F Architecture
Causal CGM provides versatile initialisation options for the adjacency matrix $A$, tailored to specific scenarios. In certain instances, weights can be derived from domain expertise or provided along with training data and labels. Without such information, weights can be directly inferred from training labels through causal structural learning algorithms (Kaddour et al., 2022) as a preliminary step. These approaches guide the model towards a predetermined decision-making pathway. Alternatively, weights can be learnt concurrently during the Causal CGM training phase, as outlined in Section 3.3. These weights may be initialised either randomly or based on the conditional entropy between labels, providing a better starting point. Additionally, a hybrid approach is feasible, where certain elements in $A$ are fixed while others remain trainable. For instance, a causal structural learning algorithm might yield a Partial Ancestral Graph (PAG) with undirected edges, allowing for the definition of directed edges and learning the direction for others to avoid cycle formation.
@@
In Causal CGM, the function $f_{i}$ corresponds to a causal mechanism in a SCM. Such mechanisms are typically formalised via structural equations. For instance, linear models are a common choice for label predictors in Concept Bottleneck Models (Koh et al., 2020) where endogenous embeddings are aggregated using a permutation invariant aggregator function $\oplus$ (such as the element-wise maximum, or sum):
$\hat{v}_{i}=\sigma\left(W_{i}\bigoplus_{j\in\{1,\dots,k\}}a_{ij}\hat{\mathbf{v}}_{j}+\mathbf{b}_{i}\right)$ (16)
However, other options are also available to increase the expressiveness and interpretability of the decision-making process, such as Deep Concept Reasoning (Barbiero et al., 2023) class predictors, which build logic-based formulae to obtain class label predictions using endogenous embeddings:
$\hat{v}_{i}\leftarrow\bigvee_{\mathbf{x}\in X_{\text{train}}}\bigwedge_{j\in\{1,\dots,k\}}l_{j}(\mathbf{x})=\bigvee_{\mathbf{x}\in X_{\text{train}}}\bigwedge_{j\in\{1,\dots,k\}}(\rho_{ij}(a_{ij}\hat{\mathbf{v}}_{j})\iff\hat{v}_{j}^{\prime})$ (17)
@@
Compositional generalisation
The training procedure of Causal CGMs is highly parallelizable and modular as only direct connections need to be trained together (e.g., $a\rightarrow b$ and $b\rightarrow c$), while the model takes care of distant connections in an indirect way. For instance, the connection $a\rightarrow b\rightarrow c$ can be obtained as a composition of two different independent training procedures for $a\rightarrow b$ and $b\rightarrow c$. As a result, it is trivial for a Causal CGM to make the causal graph grow even at test time (see Figure 11). This can be done by composing two different graphs obtained by independent training procedures, encoders, datasets, or data types. We note that this is not possible in standard CBMs, which need to re-train the task predictor from scratch whenever new concepts or tasks are added to the mix. Moreover, this modularity enables a form of out-of-distribution compositional generalisation as it creates new distant connections between variables that were never part of the same training procedure (e.g., $a$ and $c$ in the previous example).
Figure 11: Compositional generalisation in Causal CGMs: two different Causal CGMs architectures are trained independently and then composed only at test time, thus creating a larger graph and allowing out-of-distribution causal inference.
Appendix G Experimental setup
G.1 Datasets
@@
  • dSprites (Matthey et al., 2017) — The dataset comprises images featuring one of three objects (square, heart) in various positions and sizes. The defined concepts include: (1) object shape (square or heart), (2) object size (small or large), (3) vertical position (top or bottom), (4) horizontal position (left or right), (5) object colour (red or blue). Based on these, causal relationships and a binary classification task are established: if the object is a heart on the right side, it is large; if a heart is at the top of the image, it is red; the label is positive if the object is both red and large.
  • CelebA (Liu et al., 2015) — The CelebA dataset features celebrity images annotated with various attributes, including lipstick presence, gender, facial shape, and hair type. Gender is used as the classification label. This dataset is chosen for the presence of correlations and biases, such as the association between wearing lipstick and being identified as female.
  • CIFAR-10 (Krizhevsky, 2009) — The CIFAR-10 dataset consists of 60,000 colour images in 10 different classes, such as cats, dogs, and cars, with no pre-defined concepts. To extract concepts, we follow the Label-Free Concept approach described in Oikarinen et al. (2023), where the concept set is extracted using auxiliary Vision Language Models (VLS), which are then used to establish causal relationships and perform classification tasks. This dataset is used to test our model’s ability to handle complex visual features in a setting where concepts are not obtained through human annotations.
G.2 Baselines
@@
  • Concept Bottleneck Model (CBM) (Koh et al., 2020): This model first uses a concept predictor to infer concepts from the initial input, followed by a task label prediction based on these concepts. It is designed to be interpretable and treats concepts as significant informational to predict the task label.
  • Concept Embedding Model (CEM) (Espinosa Zarlenga et al., 2022): Comprising $n$ context encoders, one for each concept, this model predicts each concept based on its respective context before predicting the final task label. It treats concepts and task labels in the same way as CBM.
G.3 Experiments
In our experiments, we evaluate our approach by examining four key dimensions: (i) performance accuracy, (ii) influence of ground-truth interventions, (iii) identification of causal structures, and (iv) blocking for the influence of one variable on another.
@@ G.3 Experiments
In the second aspect, we evaluate our approach by comparing it against CBM and CEM in terms of response to ground-truth interventions. Enhancing the impact of interventions in the Concept-Based Model is crucial for improving the role of humans in the loop. In our experiments, we initially perturbed the inputs to reduce label prediction accuracy, following methodologies established in prior research (Espinosa Zarlenga et al., 2022). Subsequently, we implemented interventions on the most inaccurately predicted concepts in CEM and CBM. This intervention strategy is considered highly effective, as noted in (Shin et al., 2023). For the Causal CGM, interventions began with concepts that have a higher number of descendant nodes in the model’s graph, aiming to maximise the intervention’s effectiveness. To assess this dimension, we measured the change in accuracy for non-intervened labels before and after the interventions on $n$ concepts (Delta Label Accuracy).
In the third aspect, we visualise the DAG utilised during the inference stage by Casual CEM and derive the corresponding logic equations. We generate Sum of Product logic rules from a table that lists all possible combinations of input concept values alongside the most frequent prediction for each combination derived from the training set, similarly to what done by (Ciravegna et al., 2023). It is crucial to note that while these logic rules are general for the model decision-making process, exogenous information may alter predictions for particular instances.
In the final dimension of our analysis, we compare Causal CGM, which operates on a specified graph, against CBM and CEM in terms of their efficacy in mitigating a variable’s influence on the task. Specifically, we perturb the prediction of a concept and then, following the graph structure utilised by Causal CGM, we intervene with the ground truth label on all descendant nodes (blocking). Consequently, in Causal CGM, all links between the altered concepts and the labels are deleted, a feat unachievable in CBM and CEM. To assess this characteristic, we calculated the Residual Concept Causal Effect, ratio of the Concept Causal Effect (Goyal et al., 2020) post- and pre-application of the blocking techniques. Ideally, this ratio should be zero, indicating that after blocking, the altered node’s value no longer influences the outcome of the task.
G.4 Implementation details
Additional details To maximise the efficacy of interventions in Causal CGM, the second term of the loss can be regularised to maximise the average Causal Concept Effect (CaCE) (Goyal et al., 2020) as follows:
@@ G.4 Implementation details
Here, $r$ represents a randomly chosen index for each sample, which is used to select one of the concepts following Espinosa Zarlenga et al. (2022). This regularisation can be weighted using an hyperparameter, $\lambda_{3}$. Moreover, for all the experiments where the graph is learnt end-to-end, we initialise the learnable adjacency matrix with the conditional entropy between each pair of values, extracted from the training set.
@@
Hyperparameters All baseline and proposed models were trained for varying epochs across different datasets: 500 for Checkmark, 200 for dSprites, 30 for CelebA and 25 CIFAR10. The optimal epoch for each was determined based on label accuracy on the validation set. A uniform learning rate of 0.01 was applied across all models and datasets. For the CBM and CEM models, both concept and task losses were equally weighted at 1. This weighting scheme was also applied to the loss terms for endogenous copies’ prediction, endogenous variables’ prediction ($\lambda_{1}$), and graph priors ($\lambda_{2}$). The weight assigned to the loss terms in our models to maximize CaCE is 0.05. Additionally, $\gamma$ was treated as a learnable parameter, initialized at 0.1, and $\beta$ was set to 1. All experiments were conducted using five different seeds (1, 2, 3, 4, 5).
@@
Code, licenses and hardware For our experiments, we implement all baselines and methods in Python 3.9 and relied upon open-source libraries such as PyTorch 2.0 (Paszke et al., 2019) (BSD license), PytorchLightning v2.1.2 (Apache Licence 2.0), Sklearn 1.2 (Pedregosa et al., 2011) (BSD license). In addition, we used Matplotlib (Hunter, 2007) 3.7 (BSD license) to produce the plots shown in this paper. Two datasets we used are freely available on the web with licenses: dSprites (Apache 2.0) and CelebA, which is released for non-commercial use research purposes only. We also introduce the Checkmark dataset and we described it in this section. We will publicly release the code with all the details used to reproduce all the experiments under an MIT license. All the experiments except the CIFAR10 ones were performed on a device equipped with an M3 Max and 36GB of RAM, without the use of a GPU. The CIFAR10 experimentswere conducted on a workstation equipped with an NVIDIA RTX A6000 GPU, two AMD EPYC 7513 32-Core processors, and 512 GB of RAM.Approximately 60 hours of computational time were utilized from the start of the project, whereas reproducing the experiments detailed here requires only 8 hours.
Appendix H Additional results
@@
In this section, we include all the experiments shown in Section 4 for the four datasets in more detail and an ablation study on the value of $\lambda_{3}$.
H.1 Ground-Truth interventions
@@
Figure 12 illustrates the performance comparison among Causal CGM (using both the provided and learnt graphs), CBM, and CEM regarding the effects of interventions. Delta Label Accuracy, which quantifies the change in label accuracy before and after interventions on a growing number of concepts, is calculated solely for the concepts not directly intervened upon. In particular, Causal CGM demonstrates superior performance when interventions involve fewer concepts. This superior performance is attributed to the propagation of intervention effects through all descendant nodes in Causal CGM, unlike CBM and CEM, where the impact is confined to the final task without adjustments to other concepts. The most significant performance gain, approximately 15 percentage points, is observed in CelebA after seven interventions. This effect is particularly pronounced in scenarios with multiple concepts, such as in CIFAR10 and CelebA. In contrast, for simpler tasks with fewer concepts, like Checkmark, the benefits of our approach are less pronounced. The gap observed in CIFAR-10 between CBM and Causal CGM after all interventions is partly due to our evaluation setup: when perturbing the input prior to intervention, we occasionally generate out-of-distribution (OOD) concept embeddings, which only affected the CIFAR10 experiment. This reduces the effectiveness of interventions for embedding-based models, while scalar-based interventions in CBMs remain unaffected. Despite this challenge, our model still achieved approximately 15% higher accuracy overall, with fewer required interventions. Furthermore, when focusing solely on the effects of interventions on the task label, the causal graph utilised by Causal CGM allows us to identify beforehand the specific subset of concepts influencing the task prediction. This pre-identification significantly decreases the required number of interventions to achieve the desired outcome. Figure 13 demonstrates that Causal CGM attains comparable improvements in task performance but after interventions on only three or four concepts, in contrast to the ten and eleven concepts required by CEM and CBM, respectively. The elevated standard error observed in Causal CGM with the learnt graph is attributed to the variability of the graph structure, which significantly influences the outcomes of interventions.
H.2 Causal structures
@@
Figure 14 illustrates the adjacency matrices corresponding to the DAGs used by Causal CGM for inference in the first three datasets. On the other hand, Tables 4, 5, and 6 present the logic rules derived from the adjacency matrices depicted in the aforementioned figure. Notably, in the Checkmark dataset, both configurations successfully identified the ground truth graph and the correct logic rules. In the case of dSprites, the DAG identified through causal structural learning (GRaSP (Lam et al., 2022)) accurately discovers the causal graph and associated logic rules. Although the end-to-end model accurately identifies the correct relationships between concepts and tasks, it proposes alternative methods for concept prediction. It is important to note that even though the model did not identify the correct causal graph, the model was still capable of performing causal inference with the existing graph. In the CelebA data set, where there is no ground truth for either the graph or logic rules, the findings by GRaSP and the end-to-end model appear plausible and reveal biases inherent in the dataset, such as the strong correlation between makeup use and gender or potential causal links like smiling and a slightly open mouth. This scenario underscores the benefits of employing Causal CGM, particularly in demonstrating how specific concepts are used to predict other concepts and tasks.
H.3 Ablation study
@@
In Tables 7, 8, and 9, we present the outcomes of varying the hyperparameter $\lambda_{3}$, which weights the loss term designed to enhance the CaCE effect. The results indicate that optimising this loss term contributes to improved CaCE scores, thereby augmenting the efficacy of the interventions. Nonetheless, excessively high values of $\lambda_{3}$ may lead to diminished model performance, as it tends to prioritise boosting the CaCE score at the expense of accurate predictions.
@@
Figure 12: Impact of ground-truth interventions on concepts across four datasets. This figure illustrates the variations in accuracy for non-intervened labels, comparing performance before and after interventions on specific nodes.
@@
Figure 13: Impact on the task accuracy of ground-truth interventions performed on CelebA concepts. Causal CGM+CD has received in input a causal graph, discovered with a causal structural learning algorithm (GRaSP (Lam et al., 2022)), while Causal CGM learns it end-to-end.
@@
Figure 14: Adjaciency matrices representing the DAG used by Causal CGM during inference on the three datasets. On the left side, the matrices represent the DAG learnt end-to-end by the model, while on the right the DAG discovered with GRaSP (Lam et al., 2022). It provides a PAG starting from the training data.
@@
Table 4: Logic rules extracted for the Checkmark dataset from Causal CGM+CD with a given DAG and from Causal CGM with a learnt DAG. A term which refers to an exogenous variable is omitted for simplicity.
@@
Causal CGM $\text{a}&\leftarrow\epsilon_{0}\\ \text{b}&\leftarrow\epsilon_{1}\\ \text{c}&\leftarrow\sim\text{b}\\ \text{d}&\leftarrow\text{a}\wedge\text{c}\\ \end{aligned}$
Causal CGM+CD $\text{a}&\leftarrow\epsilon_{0}\\ \text{b}&\leftarrow\epsilon_{1}\\ \text{c}&\leftarrow\sim\text{b}\\ \text{d}&\leftarrow\text{a}\wedge\text{c}\\ \end{aligned}$
Table 5: Logic rules extracted for the Dsprites dataset from Causal CGM+CD with a given DAG and from Causal CGM with a learnt DAG. A term which refers to an exogenous variable is omitted for simplicity.
@@
Causal CGM $\text{Shape}&\leftarrow\text{Size}\\ \text{Size}&\leftarrow\epsilon_{1}\\ \text{PosY}&\leftarrow\epsilon_{2}\\ \text{PosX}&\leftarrow\epsilon_{3}\\ \text{Color}&\leftarrow\text{Shape}\\ \text{Label}&\leftarrow\text{Size}\wedge\text{Color}\\ \end{aligned}$
Causal CGM+CD $\text{Shape}&\leftarrow\epsilon_{0}\\ \text{Size}&\leftarrow\text{Shape}\wedge\text{PosX}\\ \text{PosY}&\leftarrow\epsilon_{2}\\ \text{PosX}&\leftarrow\epsilon_{3}\\ \text{Color}&\leftarrow\text{Shape}\wedge\text{PosY}\\ \text{Label}&\leftarrow\text{Size}\wedge\text{Color}\\ \end{aligned}$
Table 6: Logic rules extracted for the Celeba dataset from Causal CGM+CD with a given DAG and from Causal CGM with a learnt DAG. A term which refers to an exogenous variable is omitted for simplicity.